OpenStudy (anonymous):

perform the following operations. 37.897-3(4.07-8.252)

OpenStudy (anonymous):

41.795 use the order of operations 1.) Distribute property -3*(4.05*-8.252) -12.15+8.252=-3.898 2.) 37.897-(-3.898)=41.795

OpenStudy (anonymous):

Remember the mnemonic pemdas, please excuse my dear aunt sally,parentheses, exponet, multiplication, division, addition, subtraction

OpenStudy (anonymous):

Actually, if you consider the order of operation's acronym, P, "parenthesis" comes first, where you must do the following: = 37.897 - 3(4.07 - 8.252) = 37.897 - 3(-4.182) = 37.897 - (-12.546) = 50.443